OpenStudy (anonymous):

Will award medal! Find the most general antiderivative. ∫ (2x^3-7x+2) dx

OpenStudy (anonymous):

@Hero @Loser66

OpenStudy (anonymous):

@dumbcow @AravindG

OpenStudy (anonymous):

\[\int\limits_{}^{} x ^{n}= \frac{ x ^{n+1} }{ n+1 }\]

OpenStudy (dumbcow):

its a polynomial, these integrals are straightforward \[\int\limits x^{n} = \frac{1}{n+1} x^{n+1}\]

OpenStudy (anonymous):

Could you help me with the steps on what to do to get the answer?

OpenStudy (dumbcow):

we did...apply the formula to each term

OpenStudy (anonymous):

That's what I'm stuck on

OpenStudy (loser66):

why are you stuck?? quite easy. \[\int (2x^3 - 7x +2) dx = \int 2x^3dx-\int 7xdx +\int 2dx\] no substitute, no chain rule, no product rule, no.... nothing, hehehe, just take integral as usual. that's it.

OpenStudy (anonymous):

make sure to remember... the most GENERAL antiderivative is that with an undetermined constant \(C\) ;-)

OpenStudy (anonymous):

Thanks @Loser66 and @oldrin.bataku! :)